Calculate the net energy change in kilojoules per mole that takes place on formation of BeF2(s) from the elements: Be(s)+F2(g)⟶BeF2(s) The following information is needed:

Heat of sublimation for Be(s)= 325.8 kJ/mol

Eea for F(g)= −328 kJ/mol

Bond dissociation energy for F2(g)= 158 kJ/mol

Ei1 for Be(g)= 899.5 kJ/mol

Electrostatic interactions in BeF2(s)= −3505 kJ/mol

Ei2 for Be(g)= 1757.1 kJ/mol
Answer #1

Be(s) +52(9) BeFx (5) AE = ? Bels) —> Belo) Esublimation = 325.8*7 mont Be(g)-12> Bet (9) Eil = 899.5kJ moi? Bet(s) – sé Bétc